OpenStudy (anonymous):

1/(x^2-2)^3 differentiate with respect to x help plzzz

OpenStudy (rogue):

Use the chain rule. \[\frac {d}{dx} u^a = a u^{a-1} \frac {du}{dx}\]\[\frac {d}{dx} (x^2 - 2)^{-3} = -3(x^2 -2)^{-4} \frac {d}{dx} (x^2 - 2) = 2x * -3(x^2 -2)^{-4}\]\[\frac {d}{dx} (x^2 - 2)^{-3} = \frac {-6x}{(x^2 - 2)^4}\]